# Targeted Latent Adversarial Training Improves Robustness to Persistent Harmful Behaviors in LLMs

### v1

@@ 4.3.1 Who’s Harry Potter?
We work with the “Who’s Harry Potter” (WHP) method from Eldan and Russinovich [55]. Their method involves taking a corpus of text to forget (e.g., the Harry Potter books), constructing alternative genericized text for that corpus, and then fine-tuning the model on the generic labels. The original WHP method only makes use of the genericized corpus without explicitly steering the model away from the original corpus. Because our goal is to augment WHP with LAT, as a baseline, we use a modified version of WHP, which we call WHP-Contrastive (WHP-C). As with our SFT, R2D2, and DPO baselines from the above experiments, WHP-C trains the model with a contrastive training objective that contains both a “toward” and “away” loss. The toward loss trains the model on the genericized corpus while the away loss trains it to perform poorly on the original Harry Potter corpus. Also as before, we interleave supervised fine-tuning batches on the UltraChat dataset [130] to stabilize training when performing WHP-C. When performing WHP-C-LAT, we optimize the adversarial attacks to minimize the cross-entropy loss on the original Harry Potter text. For all methods, we train on 100 batches of size 16 for 4 steps each. Finally, in Appendix D, we also experiment with optimizing and constraining adversarial perturbations in a whitened space before de-whitening and adding them to the model’s latents.
@@
To evaluate performance on non-target tasks we again use MMLU [131]. Next, we evaluate Harry Potter familiarity [55] under atypical use cases in which an adversary may be working to extract unintended knowledge from the model. Full details are available in Appendix E. First, in response to past work suggesting that unlearning can fail to transfer cross-lingually [40], we evaluate familiarity in Spanish. Second, to test the robustness of unlearning to jailbreaks [40], we evaluate familiarity under jailbreaking prompts [141]. Third and fourth, we evaluate the extent to which the model is robust to knowledge extraction attacks [107, 111, 28, 39, 40] in the form of high-level summaries and short snippets of text from the Harry Potter books.
@@ 4.3.2 Unlearning WMDP Biology and Cyber Knowledge
As in as in Li et al. [57], we use the WMDP biology and cyber corpora as forget datasests and WikiText [142] as a retain dataset. Note that, unlike in the experiments above, the forget and retain corpora were separate and did not come as paired sets of desirable/undesirable completions for a prompt.
@@
We evaluate how well the model’s general capabilities have been preserved by testing on MMLU [131] and AGIEval [144]. We evaluate the effectiveness of unlearning in the model using biology and cyber knowledge assessments from Li et al. [57]. To test the robustness of the unlearning, we also evaluate models under few-shot finetuning attacks in which an attacker seeks to extract knowledge by finetuning the model on a small number of examples [25, 41, 42, 43, 44, 45, 46]. Here, we use a simple but surprisingly effective attack: we randomly sample a single batch of 2 examples from the relevant forget set and repeatedly train on that single batch for 20 iterations. We then report the highest WMDP bio/cyber performances for each model across evaluation checkpoints at 5, 10, and 20 steps. For all evaluations, we use 1,000 samples on lm-evaluation-harness v0.4.0 [145] as done in Li et al. [57].
@@ 5 Discussion
By attacking the model’s latent representations, LAT offers a unique solution because models represent concepts at a higher level of abstraction in the latent space [2]. Here, we have used targeted latent adversarial training (LAT) to strengthen existing defenses against persistent harmful behaviors in LLMs. We have tested targeted LAT by applying it to three current challenges with state-of-the-art LLMs: jailbreaking [21], unlearning [105], and backdoor removal [103, 102]. In each case, we have shown that LAT can augment existing techniques to improve the removal to unwanted behaviors with little or no tradeoff in general performance. This is especially notable in the case of backdoor removal in which DPO alone fails to remove backdoors while DPO with LAT does so very effectively. This suggests that LAT can be a solution to the “Sleeper Agent” problem posed in Hubinger et al. [54] and Pawelczyk et al. [98] in which harmful backdoors can persist through adversarial training.
@@
Our motivation for LAT is a response to two observations. First, LLMs empirically can persistently retain harmful capabilities despite attempts to remove them with adversarial training [3, 20, 25, 29, 30, 41, 42, 43, 44, 45, 46, 126, 141]. Second, there have been empirical and theoretical findings that LLMs undergo limited changes to their inner capabilities during fine-tuning [24, 25, 26, 27, 33, 34, 35, 36, 37, 38, 39]. All three problems that we have used targeted LAT to address – jailbreaks, backdoors, and undesirable knowledge – are ones in which an LLM exhibits harmful behaviors that are difficult to thoroughly remove. Our results show that targeted LAT can be useful for making models more robust to known failures in high-stakes settings. In practice, we envision LAT being useful as an effective defense alongside other methods to filter harmful model inputs and outputs [83, 146, 147, 148].
@@
In Section 4.3, we demonstrate that state-of-the-art LLM unlearning methods are surprisingly vulnerable to relearning from small amounts of data. Even with state-of-the-art methods, we find that re-training repeatedly on only two samples from the forget set was able to close more than half of the performance gap between the original and unlearned models on average. We find that targeted LAT can reduce the sample efficiency of re-learning, but there remains much room for improvement in designing unlearning methods that are robust to few-shot finetuning attacks. We are interested in future work to explore LAT’s potential to improve on existing approaches for making models robust to few-shot fine-tuning attacks [149, 150, 151, 152].
@@
Improved latent-space attacks In addition to performing LAT with perturbations to an LLM’s residual stream, we are interested in other strategies for attacking its internal representations. Toward this goal, engaging with recent work on LLM representation engineering [2, 82] and interpretability [153] may help to parameterize and shape latent space attacks. We also speculate that universal attacks instead of single-instance attacks may be more interpretable and might better target the most prominent mechanisms that a model uses when it produces undesirable outputs.
Augmenting circuit breaking Concurrently with our work, Zou et al. [154] introduced a “circuit breaking” method for significantly improving LLM robustness to jailbreaks. We are interested in augmenting circuit breaking with LAT to further improve robustness.
@@ Acknowledgements
We are thankful to Rajashree Agarwal, Rohit Gandikota, John Hughes, Erik Jenner, Alex Lyzhov, Sam Marks, Jacob Pfau, Sara Price, Javier Rando, Markian Rybchuk, Lennart Schulze, Aaquib Syed, Alex Turner, and Tony Wang for useful conversations. We thank William Brewer, Rocket Drew, Ronny Fernandez, McKenna Fitzgerald, Juan Gil, Carson Jones, Ryan Kidd, Christian Smith, and Laura Vaughan, for program support. This project was funded in part by a grant from Open Philanthropy and used compute provided by the Center for AI Safety.
@@ A.1 R2D2-LAT
Note that is calculated on inputs where no adversarial attack is present.
@@ A.4 RMU-LAT
where is the length of the input tokens, and u is a randomly chosen vector from a uniform distribution between that is then normalized (and stays constant throughout training). The constants and are hyperparameter coefficients, which we set to be 6.5 and 1200 as referenced in Li et al. [57] for Zephyr-7B.
Appendix B Jailbreaking Robustness Under an Alternate Autograder
@@
Model General Performance Attack Success Rate Relative MMLU MT-Bench Compliance Direct Req. Prefill AutoPrompt GCG Many-Shot Compute Llama2-7B-chat 0.464 0.633 0.976 0.000 0.594 0.229 0.417 0.949 0x RT 0.456±0.012 0.632±0.045 0.936±0.035 0.049±0.027 0.226±0.096 0.285±0.144 0.490±0.240 0.458±0.181 1x R2D2 0.441±0.001 0.569±0.029 0.938±0.021 0.000±0.000 0.215±0.021 0.007±0.003 0.028±0.007 0.111±0.003 6558x RT-EAT 0.448±0.003 0.622±0.002 0.944±0.028 0.010±0.000 0.146±0.095 0.021±0.000 0.080±0.013 0.000±0.000 9x RT-EAT-LAT (ours) 0.454±0.001 0.586±0.007 0.962±0.016 0.003±0.003 0.122±0.048 0.021±0.004 0.018±0.007 0.000±0.000 9x Llama3-8B-chat 0.638 0.839 1.000 0.104 0.729 0.271 0.596 0.323 0x RT 0.631 0.816 1.0 0.104 0.406 0.229 0.552 0.094 1x RT-EAT-LAT (ours) 0.612 0.787 0.986 0.000 0.177 0.000 0.021 0.000 9x
@@
Appendix C Backdoored Model MMLU Performance
Appendix D Low Rank Adapters and Scaled Perturbation Constraints for WHP Unlearning
@@
Appendix E Tests for Robust and Competitive Unlearning in LLMs
@@
Appendix F WMDP Unlearning Details
@@
There are three layer choices that can be varied in our setup: which layer(s) of the model to put the adversary, which layers to train for RMU, and which layer to do the RMU MSE activation matching over. We kept to the layers (trainable and RMU matching) for RMU stated in Li et al. [57] – the RMU layer for the activation matching, with trainable to keep the set of hyperparameters to search over reasonably small. Applying attacks to layer requires a smaller ball radius for our random perturbations; else, we found that the adversary prevents the model trained with RMU from successfully unlearning. We also find the greatest benefit in applying attacks to the layer before the RMU activation matching layer.

### v2

@@ 4.3.1 Who’s Harry Potter?
We work with the “Who’s Harry Potter” (WHP) method from Eldan and Russinovich [56]. It involves taking a corpus of text to forget (e.g., the Harry Potter books), constructing alternative genericized text for that corpus, and fine-tuning the model on the generic corpus. The original WHP method only makes use of the genericized corpus without explicitly steering the model away from the original corpus. Because our goal is to augment WHP with LAT, as a baseline, we use a modified version of WHP, which we call WHP-Contrastive (WHP-C). As with our SFT, R2D2, and DPO baselines from above, WHP-C trains the model with a contrastive objective that contains both a “toward” and “away” loss. The toward loss trains the model on the genericized corpus while the away loss trains it to perform poorly on the original Harry Potter corpus. Also as before, we interleave supervised fine-tuning batches on the UltraChat dataset [133] to stabilize training. When performing WHP-C-LAT, we optimize the adversarial attacks to minimize the cross-entropy loss on the original Harry Potter text. For all methods, we train on 100 batches of size 16 for 4 steps each. Finally, in Appendix E, we also experiment with optimizing and constraining adversarial perturbations in a whitened space before de-whitening and adding them to the model’s latents.
@@
To evaluate general performance, we again use MMLU [134]. Next, we evaluate Harry Potter familiarity [56] under Harry Potter knowledge extraction attacks. Full details are available in Appendix F. First, in response to past work suggesting that unlearning can fail to transfer cross-lingually [40], we evaluate familiarity in Spanish. Second, to test the robustness of unlearning to jailbreaks [40], we evaluate familiarity under jailbreaking prompts [144]. Third and fourth, we evaluate the extent to which the model is robust to knowledge extraction attacks [110, 114, 28, 39, 40] in the form of high-level summaries and short snippets of text from the Harry Potter books.
@@ 4.3.2 Unlearning WMDP Biology and Cyber Knowledge
We evaluate how well the model’s general capabilities have been preserved by testing on MMLU [134] and AGIEval [147]. We evaluate the effectiveness of unlearning in the model using biology and cyber knowledge assessments from Li et al. [58]. These multiple choice evaluations represent a qualitatively different task than the forget sets (which were full of bio and cyber documents), so they test the ability of LAT to generalize to qualitatively different kinds of unwanted behaviors than those used during fine-tuning. To test the robustness of the unlearning, we also evaluate models under few-shot finetuning attacks in which an attacker seeks to extract knowledge by finetuning the model on a small number of examples [25, 41, 42, 43, 44, 45, 46, 50]. Here, we use a simple but surprisingly effective attack: we randomly sample a single batch of 2 examples from the relevant forget set and repeatedly train on that single batch for 20 iterations. We then report the highest WMDP bio/cyber performances for each model across evaluation checkpoints at 5, 10, and 20 steps. For all evaluations, we use 1,000 samples on lm-evaluation-harness v0.4.0 [148] as done in Li et al. [58].
@@ 5 Discussion
By attacking the model’s latent representations, LAT offers a unique solution because models represent concepts at a higher level of abstraction in the latent space [2]. Here, we have used targeted latent adversarial training (LAT) to strengthen existing defenses against persistent harmful behaviors in LLMs. We have applied LAT to three current challenges with state-of-the-art LLMs: jailbreaking [21], unlearning [108], and backdoor removal [106, 105]. In each case, we have shown that LAT can augment existing techniques to improve the removal of unwanted behaviors with little or no tradeoff in general performance. Overall, these results support but do not yet confirm our hypothesis that LAT can remove neural circuitry from models responsible for undesirable behaviors. We leave analysis of the mechanisms behind harmful model behaviors (e.g., [85]) to future work.
@@
Our motivation for LAT is a response to two observations. First, LLMs empirically can persistently retain harmful capabilities despite attempts to remove them with adversarial training [3, 20, 25, 29, 30, 41, 42, 43, 44, 45, 46, 129, 144]. Second, there have been empirical and theoretical findings that LLMs undergo limited changes to their inner capabilities during fine-tuning [24, 25, 26, 27, 33, 34, 35, 36, 37, 38, 39]. All three problems that we have used targeted LAT to address – jailbreaks, backdoors, and undesirable knowledge – are ones in which an LLM exhibits harmful behaviors that are difficult to thoroughly remove. Our results show that targeted LAT can be useful for making models more robust to these persistent failures. We also find that these failure modes need not be precisely known for LAT to be helpful, showing instances in which LAT can improve generalization to different datasets of attack targets, harmful behaviors, and knowledge-elicitation methods than were used during training.
@@
In Section 4.3, we find that state-of-the-art LLM unlearning methods are surprisingly vulnerable to relearning from small amounts of data. We find that re-training repeatedly on only two samples from the forget set was consistently able to close more than half of the performance gap between the original and unlearned models on average. We find that targeted LAT can reduce the sample efficiency of re-learning, but there is much room for improvement in designing unlearning methods that are robust to few-shot finetuning attacks. We are interested in future work to explore LAT’s potential to improve on existing approaches for making models robust to few-shot fine-tuning attacks [149, 150, 151, 152].
@@
Improved latent-space attacks In addition to performing LAT with perturbations to an LLM’s residual stream, we are interested in other strategies for attacking its internal representations. Toward this goal, engaging with recent work on LLM representation engineering [2, 84] and interpretability [153] may help to better parameterize and shape latent space attacks. We also speculate that universal attacks instead of single-instance attacks may be more interpretable and might better target the most prominent mechanisms that a model uses when it produces undesirable outputs.
Augmenting other latent-space manipulation techniques Concurrently with our work, Zou et al. [154], Rosati et al. [152], and [155] introduced other latent-space manipulation techniques for making LLMs robust to undesirable behaviors. We are interested in studying how these techniques compare to LAT and whether LAT can be used to improve them.
@@ Acknowledgements
We are thankful to Rajashree Agarwal, Rohit Gandikota, John Hughes, Erik Jenner, Alex Lyzhov, Sam Marks, Jacob Pfau, Sara Price, Javier Rando, Markian Rybchuk, Lennart Schulze, Aaquib Syed, Alex Turner, and Tony Wang for useful conversations. We thank William Brewer, Rocket Drew, Ronny Fernandez, McKenna Fitzgerald, Juan Gil, Carson Jones, Ryan Kidd, Christian Smith, and Laura Vaughan, for program support. This project was funded in part by a grant from Open Philanthropy and used compute provided by the Center for AI Safety. This support was offered by both organizations without conditions, and neither exerted any influence over the paper’s content.
@@ A.1 RT-LAT
We can use one of two different benign loss terms:
where are the weights of the frozen reference model. Note that is always calculated on inputs where no adversarial attack is present.
We use for our Llama2 results, and for our Llama3 experiments. trains the model to maximize the probability of the ground-truth completions for benign prompts, whereas trains the model to preserve its original logits over possible completions for benign prompts. We hypothesize that might preserve original model capabilities better when the quality of is poor relative to the model being trained. Empirically, we find that can better allow more capable models to retain their capabilities during adversarial training.
@@ A.4 RMU-LAT
where is the length of the input tokens, and u is a randomly chosen vector from a uniform distribution between that is then normalized (and stays constant throughout training). The constants and are hyperparameter coefficients, which we set to be 6.5 and 1200 as in Li et al. [58] for Zephyr-7B.
Appendix B Jailbreaking Robustness Under Untargeted LAT
To test the advantages of targeted LAT over untargeted LAT, we compare the jailbreaking robustness of the two in Table 6. Here, during untargeted LAT, the adversary does not work to make the model comply with the jailbreak. Instead, it only works to make the model fail to output a refusal. We find that untargeted LAT results in less harm to general performance compared to targeted LAT but not refusal training. Meanwhile, untargeted lat results in comparable or slightly worse robustness in most cases compared to targeted LAT. However, for prefill and GCG attacks, untargeted LAT fares much worse than targeted LAT.
Model General Performance Attack Success Rate Relative MMLU MT-Bench Compliance Direct Req. PAIR Prefill AutoPrompt GCG Many-Shot Compute Llama3-8B-instruct 0.638 0.839 1.000 0.086 0.089 0.488 0.151 0.197 0.165 0x RT 0.639±0.000 0.836±0.009 1.000±0.000 0.000±0.000 0.143±0.010 0.135±0.016 0.010±0.004 0.039±0.012 0.033±0.009 1x RT-EAT-LAT (untargeted) 0.636±0.001 0.836±0.004 0.999±0.001 0.000±0.000 0.099±0.003 0.375±0.013 0.007±0.004 0.076±0.004 0.000±0.000 9x RT-EAT-LAT (ours) 0.613±0.009 0.829±0.013 0.998±0.000 0.000±0.000 0.033±0.010 0.068±0.021 0.000±0.000 0.009±0.002 0.000±0.000 9x
Appendix C Jailbreaking Robustness Under an Alternate Autograder
@@
Model General Performance Attack Success Rate Relative MMLU MT-Bench Compliance Direct Req. PAIR Prefill AutoPrompt GCG Many-Shot Compute Llama2-7B-chat 0.464 0.633 0.976 0.000 0.390±0.000 0.594 0.229 0.417 0.949 0x RT 0.456±0.012 0.632±0.045 0.936±0.035 0.049±0.027 0.317±0.024 0.226±0.096 0.285±0.144 0.490±0.240 0.458±0.181 1x R2D2 0.441±0.001 0.569±0.029 0.938±0.021 0.000±0.000 0.180±0.007 0.215±0.021 0.007±0.003 0.028±0.007 0.111±0.003 6558x RT-EAT 0.448±0.003 0.622±0.002 0.944±0.028 0.010±0.000 0.177±0.008 0.146±0.095 0.021±0.000 0.080±0.013 0.000±0.000 9x RT-EAT-LAT (ours) 0.454±0.001 0.586±0.007 0.962±0.016 0.003±0.003 0.050±0.002 0.122±0.048 0.021±0.004 0.018±0.007 0.000±0.000 9x Llama3-8B-Instruct 0.638 0.839 1.000 0.104 0.540 0.729 0.271 0.596 0.323 0x RT 0.639±0.000 0.836±0.015 1.000±0.000 0.000±0.000 0.603±0.003 0.229±0.021 0.021±0.000 0.083±0.048 0.149±0.047 1x RT-EAT-LAT (ours) 0.613±0.016 0.829±0.022 0.998±0.000 0.000±0.000 0.093±0.002 0.101±0.069 0.003±0.006 0.021±0.000 0.000±0.000 9x
@@
Appendix D Backdoored Model MMLU Performance
Appendix E Low Rank Adapters and Scaled Perturbation Constraints for WHP Unlearning
@@
Appendix F Tests for Robust and Competitive Unlearning in LLMs
@@
Appendix G WMDP Unlearning Details
@@
There are three layer choices that can be varied in our setup: which layer(s) of the model to put the adversary, which layers to train for RMU, and which layer to do the RMU MSE activation matching over. We kept to the same layers (trainable and RMU matching) for RMU as in Li et al. [58] – the RMU layer for the activation matching, with trainable to keep the set of hyperparameters to search over reasonably small. Applying attacks to layer requires a smaller ball radius for our random perturbations; else, we found that the adversary prevents the model trained with RMU from successfully unlearning. We also find the greatest benefit in applying attacks to the layer before the RMU activation matching layer.